# ToDo (#1): #2

### v3

@@ 3.8.7 RF Amplifiers
We plan to use vacuum tubes (triodes or tetrodes) to produce the RF power. We have considered using LDMOS amplifiers given the recent improvements in power density, reliability, and cost. However, we believe that vacuum tubes will perform better in the harsh environment and will be able to better accommodate the rapid variations in load due to the beam temporal patterns.
@@ 3.8.8 RF Lines and Couplers
Both of these systems are capable of providing the necessary 300 kW power to each cavity. Capacitive coupling is less prone to multipacting because the coupling antenna could be easily DC biased. But, our expertise in high power coupler design advocates for an inductive solution. The design of inductive coupler is easier to water-cool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6”1/8 port. The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling. See figure 3.39. The adaptation tuning will be performed by rotation of the loop around the vertical axis. The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (figure 3.40) and a $S_{1,1}$ computed scattering parameter inferior to -60 dB.
Figure 3.39: RF loop coupler design
Figure 3.40: Coupler port impedance
@@ 3.9 Cyclotron Particle Simulations
It is necessary to perform an in-depth particle simulation of the primary acceleration of the beam to ensure that the beam reaches energy, maintains quality, and is possible to extract. We do this using OPAL  [adelmann_opal_2019] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects. It is based on particle-in-cell (PIC) calculations that track a ’bunch’ of particles through time-varying magnetic and electric fields. The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method.There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons. This makes it well-suited for the primary acceleration phase of the IsoDAR cyclotron. Conveniently, OPAL-cycl has built-in objects called PROBEs, that can provide beam diagnostics, and and COLLIMATORs, which can provide halo reduction.
@@ 3.9.1 Fieldmaps
A critical aspect of these simulations is that they accurately simulate the electric and magnetic fields within the cyclotron. We performed these simulations using a 1D RF model derived from the voltages provided in Table 3.11. The radially increasing voltage leads to a roughly constant turn separation, which is important for ease of extraction. We conducted these simulations using a 2D median plane magnetic field that is derived from a physically realistic cyclotron model produced by IBA. Of note, this magnetic field map does not include a $\nu_{r}=1$ resonance, which would increase the inter-turn separation in the last few turns at the price of reduced beam quality. This magnetic field is significantly more realistic than our previous work because the cyclotron model allots space for the injection components.
@@ 3.9.2 Probes
Probes are a built-in OPAL object that records whenever a particle passes through the region defined by the probe. These particles are simply recorded, and the trajectories of the particles are not affected by their intersection with the probe. This is helpful for determining the “extractability” of the beam. The probe collects information about the position and the momenta of the passing particles. By placing a probe at an azimuth of 25 degrees, the angle at which our electrostatic channel starts, we are able to determine the overlap between sequential turns of the beam path. This is necessary because we want to minimize this overlap for ease of extraction. The overlap between beams provides a lower bound for the power that will be deposited on the electrostatic septum, so we aim to minimize this value. The data from the probe is displayed in Fig 3.42.
@@ 3.9.3 Simulation Results
Using the field maps as discussed earlier and the set of 15 collimators as described in the previous sections. We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig 3.43. Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig 3.41. Additionally, a probe at an azimuth of 25 degrees sets a lower limit for extraction at 120W as seen in Fig 3.42.
Figure 3.41: The collimated beam had a much reduced radial RMS spread.
Figure 3.42: There are two potential extraction points
Figure 3.43: There is a smooth acceleration of the beam
@@ 3.9.4 Electrostatic Extraction Channels
The first step in our extraction is the construction of a radially directed electric field that peels the last turn away from the second to last turn. This is done by placing a series of thin (0.2 mm) grounded electrodes between the last two turns and a larger puller electrode that is at high negative potential outside of the last turn. This generates a well-shielded electric field in the region between the puller and septum. A rendering of the electrostatic septum is shown in Fig 3.44.There are two such electrostatic channels that increase separation between the last two turns.
Figure 3.44: CAD rendering of electrostatic septum. From Ref. [winklehner_new_2022]
We determine the exact location of each element of the extraction channel iteratively. The first step is to run a Python script on a pre-existing OPAL simulation of the last three turns that determines the initial positions of the septum and the puller. We choose to place the septum evenly between the last two turns. A .txt file is generated that allows for a model of the septum to be built in Autodesk Inventor. With this geometry, we use COMSOL to determine the electric field and add it to the pre-existing RF field of the cyclotron in OPAL. We model the physical dimensions of the septum as COLLIMATOR objects in OPAL. This process is repeated twice at which point further iteration does not result in any improvements in ESC performance. Following this we were able to reduce our losses on the first septum to 205 W which is 2.5% greater than our 200 W goal and we lose 21 W on the second septum. In the future we plan to reintroduce a resonance into the magnetic field which will increase the inter-turn separation at extraction which will greatly reduce this number. The end result of the septum optimization process is shown in Fig 3.45.
Figure 3.45: This is the tracking of 100 particles as they pass through the electrostatic channels (shown in green). It also shows the location of our magnetic extraction channel in grey.
@@ 3.9.5 Magnetic Extraction Channels
After the electrostatic extraction channels, the beam will rapidly de-focus as it passes through the radially decreasing magnetic field. In order to avoid this issue, we must include focusing elements that can counteract the effects of the fringe fields of the cyclotron. Due to space constraints, it is not possible to have an active element such as a quadrupole within the cyclotron. Instead, we will use the standard strategy of machining and placing iron bars along the extracted beam path. These bars allow us to locally reshape the magnetic field and are very compact. We model our magnetic extraction off of the MSU cyclotron [MSU_Cyclotron] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC. By using the gradients that are produced at MSU as a reference point, we conservatively set a gradient of 1.5 kG/cm which is much less than the 5 kG/cm that they were able to achieve. We directly modified our magnetic field to contain these elements. It is visible in Fig 3.46 that a fairly week magnetic channel is sufficient to maintain beam quality. We do not anticipate difficulties with these channels because the electrostatic channels generate the majority of the turn separation.
Figure 3.46: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
@@ 3.9.6 Risks and Mitigation
Mitigation: Understanding beam dynamics sufficiently well that this does not happen, and having adequate beam diagnostics to be able to tune the beam for maximum turn separation. Further, we could place a narrow carbon foil in front of teh septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum. The resulting protons have lower magnetic rigidity and bend away from the septum before hitting it. They follow a specific trajectory that can guide them outside the cyclotron to be dumped in a controlled way.
@@
Chapter 4 Installation in the Yemilab Setting
The most likely deployment location for IsoDAR is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea. Fig. 4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 meter vertical shaft. Unfortunately, the elevator is suitable only for personnel access. All heavy and bulky equipment must be brought down via the ramp. Fig. 4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter). The following section provides a general description of deployment strategies in an underground setting, pertaining only to the cyclotron. In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at any alternative installation sites. A similar approach will be taken for the target system in Volume II of this PDR.
Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure. From Ref. [kimYemilabNewUnderground2024].
Figure 4.2: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
@@ 4.1 Transport of Components
All of the components of the experiment must be brought to the site from the port-of-entry into the Republic of Korea. The preferred method is to use trucks along the highways. Prior experience of IBA with rail transport has been that it is difficult to provide adequate shock protection, and heavy pieces can be damaged in transit. As stated above, transport from the surface to the experimental site must be accomplished by way of the mine ramp depicted in Fig. 4.1. Nominally, its cross section is 5 x 5 meters, however we are told that pinch points and sharp bends may occur and might provide challenging areas to transport large, heavy loads through. We have received assurance that large mining trucks, 3.5 meters wide, can traverse the entire ramp, as this is the way ore was transported to the surface for many years.
@@
While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf. Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions. Improved scans must be performed as part of final installation planning in the Technical Design Report phase (TDR).
@@
As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site. The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig. 4.3 for bringing the magnet and vacuum liner pieces from the surface. These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [SPMT].
@@
Figure 4.3: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
@@
The only challenge will be the magnet coils. Each of the two coil packages measures 5 meters across and 0.3 meters high, the individual coils inside measure 4.960 x 0.252 meters. Figs. 4.5 and 4.5 gives an approximate view of a coil package and the clearances in the access ramp to the Laboratory (marked as the ”Entrance Tunnel” in Fig. 4.1 (a). This particular ramp is well-constructed and engineered, and has a measured width of 5 meters. A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through this particular passageway. There are constrictions even in this ramp: cable trays and service boxes, however these can be moved if necessary. Uncertainty remains as to whether this type of conveyance device for the coils can navigate the winding ramp from the surface.
@@
In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections. This has already been studied in Chapter 3 of the CDR document [alonso_isodaryemilab_2021, alonso_isodaryemilab_2022]. The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site. The biggest challenge would be the ventilation scheme for the potting process. There is precedence for sectioned coils. For the large TRIUMF cyclotron [TRIUMF], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways. The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
@@
Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
Figure 4.5: This ramp is engineered to a 5-meter width, and maximum height of 5 meters. Unfortunately we do not have sufficient information about the winding access ramp from the surface to the base of the shaft to assess at this time whether transport of the coils will be possible with this conveyance device.
@@ 4.2 Assembly underground
The cyclotron will be assembled in a hall measuring 12 meters wide, 17 meters deep and 10 meters high. This hall was excavated to our specifications at the 90 degree bend in the access ramp to the LSC, as seen in Fig. 4.2. A flat area, roughly 10 m x 5 m at the corner will be the staging area for components transported from the surface. A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point. Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A. [GHcranes] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig. 4.6. It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons. From Ref. [GHcranes].
Figure 4.7: Cyclotron and crane on cavern floor.
Figure 4.8: Cyclotron and crane inside the cavern.
An overview of the cyclotron inside the cavern with the rigging equipment is shown in Figures 4.8 - 4.9. It should be noted that the gantry crane will be dismantled once the heavy parts are rigged. This would increase the clearance above the cyclotron for the additional equipment. In addition, the large crane can be reconfigured and reassembled in the target hall, and used for assembly of the MEBT, target and shielding, and for subsequent maintenance of the target systems.
Figure 4.9: Cyclotron open and crane with cavern mesh - cut-view.
4.2.1 Risks and Mitigations
Risk: Coil Size. The cyclotron coil will likely not fit through all of the sections of the main access ramp at Yemilab in horizontal orientation. A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through diagonally. However, there is some uncertainty about possible choke points in the main access ramp to the Yemilab facility.
Mitigation: In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above. Both come with drawbacks or additional cost, but are manageable.
Chapter 5 Conclusion

### v4

@@ 5.9.7 RF Amplifiers
We plan to use vacuum tubes (triodes or tetrodes) to produce the RF power. We have considered using LDMOS (laterally diffused metal-oxide semiconductor) amplifiers given the recent improvements in power density, reliability, and cost. However, considering the harsh environment in question, vacuum tubes will perform more reliably and will also be better able to accommodate the rapid variations in load due to the beam temporal patterns.
@@ 5.9.8 RF Lines and Couplers
Both of these systems are capable of providing the necessary 300 kW power to each cavity. Capacitive coupling is less prone to multipactoring because the coupling antenna could be easily DC biased. But, our expertise in high power coupler design advocates for an inductive solution. The design of inductive coupler is easier to water-cool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6 1/8” port. The suggested design is a loop made out of a $20\times 60$ mm rectangular section bar, offering room for water-cooling. See Fig. 57. The adaptation tuning will be performed by rotation of the loop around the vertical axis. The design has been optimized with a default angle of 45 to the cavity axis, in order to allow a maximal tuning range.
The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (see Fig. 58) and a $S_{11}$ computed scattering parameter smaller than -60 dB.
Figure 57: CAD rendering of the RF cavity showing the position and size of the loop coupler.
Figure 58: RF loop coupler reflected power ($S_{11}$) showing excellent coupling at -60 dB.
@@ 5.10 Cyclotron Particle Simulations
It is necessary to perform an in-depth particle simulation of the primary acceleration of the beam to ensure that the beam reaches energy, maintains quality, and is possible to extract. We do this using OPAL  [55] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects. It is based on particle-in-cell (PIC) calculations that track a “bunch” of particles through time-varying magnetic and electric fields. The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method. There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons. This makes it well-suited for the primary acceleration phase of the IsoDAR cyclotron. Conveniently, OPAL-cycl has built-in objects called probes, that can provide beam diagnostics, and collimators, which can provide halo reduction.
@@ 5.10.1 Fieldmaps
A critical aspect of these simulations is that they accurately simulate the electric and magnetic fields within the cyclotron. We performed these simulations using the latest updates to our 1D RF model discussed previously. Specifically, the field is derived from the voltages provided in Table 17. The radially increasing voltage leads to a roughly constant turn separation, which is important for ease of extraction. We conducted these simulations using a 2D median plane magnetic field that is derived from a physically realistic cyclotron model produced by IBA. Of note, this magnetic field map does not include a $\nu_{r}=1$ resonance, which would increase the inter-turn separation in the last few turns at the price of reduced beam quality. This magnetic field is significantly more realistic than our previous work because the cyclotron model allots space for the injection components.
@@ 5.10.2 Probes
Probes are a built-in OPAL object that records whenever a particle passes through the region defined by the probe. These particles are simply recorded, and the trajectories of the particles are not affected by their intersection with the probe. This is helpful for determining the “extractability” of the beam. The probe collects information about the position and the momenta of the passing particles. By placing a probe at an azimuth of 25 degrees, the angle at which our electrostatic channel starts, we are able to determine the overlap between sequential turns of the beam path. This is necessary because we want to minimize this overlap for ease of extraction. The overlap between beams provides a lower bound for the power that will be deposited on the electrostatic septum, so we aim to minimize this value. The data from the probe is displayed in Fig. 60.
@@ 5.10.3 Simulation Results
Using the field maps as discussed earlier and the set of 15 collimators as described in the previous sections. We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig. 61. Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig. 59. Additionally, a probe at an azimuth of 25 sets a lower limit for extraction at 127 W as seen in Fig. 60.
Figure 59: Rms beam size in the radial direction as a function of beam energy. We observe a much reduced radial spread in the collimated beam.
Figure 60: Beam power versus radius on a radial probe inserted at 25 azimuth, binned in 0.5 mm wide bins to account for the thickness of the electrostatic septum. The final two turns exhibit low enough inter-turn power to extract the beam.
[…]
@@ 5.10.4 Electrostatic Extraction Channels
The first step in our extraction is the construction of a radially directed electric field that peels the last turn away from the second to last turn. This is done by placing a series of thin (0.2 mm) grounded electrodes between the last two turns and a larger puller electrode that is at high negative potential outside of the last turn. This generates a well-shielded electric field in the region between the puller and septum. A rendering of the electrostatic septum is shown in Fig. 62. Two such electrostatic channels increase separation between the last two turns.
Figure 62: CAD rendering of electrostatic septum. Inventor model (left). Visualization of electric field from COMSOL (right). From Ref. [6].
We determine the exact location of each element of the extraction channel iteratively. The first step is to run a Python script on a pre-existing OPAL simulation of the last three turns that determines the initial positions of the septum and the puller. We chose to place the septum evenly between the last two turns. A text file is generated that allows for a model of the septum to be built in Autodesk Inventor. With this geometry, we use COMSOL to determine the electric field and add it to the pre-existing RF field of the cyclotron in OPAL. We model the physical dimensions of the septum as COLLIMATOR objects in OPAL. This process is repeated twice at which point further iteration does not result in any improvements in ESC performance. Following this we were able to reduce our losses on the first septum to 158 W which is below our 200 W goal and we lose 7 W on the second septum. In the future, we are considering reintroducing a resonance into the magnetic field which would increase the inter-turn separation at extraction and greatly reduce this number. The end result of the septum optimization process is shown in Fig. 63.
Figure 63: 100 randomly selected particle trajectories out of the simulation sample of 200,000 particles as they pass through the electrostatic channels (shown in green). Also shown is the location of the first magnetic extraction element in gray.
@@ 5.10.5 Magnetic Extraction Channels
After the electrostatic extraction channels, the beam will rapidly de-focus as it passes through the radially decreasing magnetic field. In order to avoid this issue, we must include focusing elements that can counteract the effects of the fringe fields of the cyclotron. Due to space constraints, it is not possible to have an active element such as a quadrupole within the cyclotron. Instead, we will use the standard strategy of machining and placing iron bars along the extracted beam path. These bars allow us to locally reshape the magnetic field and are very compact. We model our magnetic extraction off of the MSU cyclotron [68] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC. By using the gradients that are produced at MSU as a reference point, we conservatively set a gradient of 1.5 kG/cm which is much less than the 5 kG/cm that they were able to achieve. We directly modified our magnetic field to mimic these elements. It is visible in Fig. 64 that a single relatively weak magnetic channel can improve beam quality. We do not anticipate difficulties implementing a series of such channels to guide and focus our extracted beam because we have significantly increased turn-to-turn separation following the electrostatic channels.
Figure 64: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
@@ 5.10.6 Risks and Mitigation
Mitigation: Understanding beam dynamics sufficiently well that this does not happen, and having adequate beam diagnostics to be able to tune the beam for maximum turn separation. Further, we could place a narrow carbon foil in front of the septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum. The resulting protons have lower magnetic rigidity and bend away from the septum before hitting it. They follow a specific trajectory that can guide them outside the cyclotron to be dumped in a controlled way.
@@
6 Installation in the Yemilab Setting
The most likely deployment location for IsoDAR is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea. Fig. 65 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 m at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 m vertical shaft. Unfortunately, the elevator is suitable only for personnel access. All heavy and bulky equipment must be brought down via the ramp. Fig. 66 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kt LSC (Liquid Scintillator Counter). The following section provides a general description of deployment strategies in an underground setting, pertaining only to the cyclotron. In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at any alternative installation sites. A similar approach will be taken for the target system in Volume II of this PDR.
Figure 65: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure. From Ref. [9].
Figure 66: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
@@ 6.1 Transport of Components
All of the components of the experiment must be brought to the site from the port-of-entry into the Republic of Korea. The preferred method is to use trucks along the highways. IBA has had prior experience with rail transport and has determined that it is difficult to provide adequate shock protection, and heavy pieces can be damaged in transit. As stated above, transport from the surface to the experimental site must be accomplished by way of the mine ramp depicted in Fig. 65. Nominally, its cross section is $5\times 5$ m, however we are told that pinch points and sharp bends may occur and might provide challenging areas to transport large, heavy loads through. We have received assurance that large mining trucks, 3.5 m wide, can traverse the entire ramp, as this is the way ore was transported to the surface for many years.
@@
While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf. Figs. 72 and 73), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions. Improved scans must be performed as part of final installation planning in the Technical Design Report phase (TDR).
@@
As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site. The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig. 67 for bringing the magnet pieces from the surface. These electrically-powered vehicles are flexible in their size, comprising sets of 4 tires, each set independently steerable, the vehicle is extremely maneuverable and can easily accommodate the heaviest piece of magnet steel [69]. The vacuum liner is made mostly of stainless steel pieces, but though large and bulky, the pieces can be brought down and welded in the underground assembly cavern.
@@
Figure 67: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
@@
The only challenge will be the magnet coils. Each of the two coil packages measures 5 meters across and 0.3 m high, the individual coils inside measure $4.960\times 0.252$ m. Figures 68 and 69 give an approximate view of a coil package and the clearances in the access ramp to the Laboratory (marked as the ”Entrance Tunnel” in Fig. 65 (a). This particular ramp is well-constructed and engineered, and has a measured width of 5 m. A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through this particular passageway. There are constrictions even in this ramp: cable trays and service boxes, however these can be moved if necessary. Uncertainty remains as to whether this type of conveyance device for the coils can navigate the winding ramp from the surface.
@@
In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections. This has already been studied in chapter 3 of the CDR document [70, 1]. The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site. The biggest challenge would be the ventilation scheme for the potting process. There is precedence for sectioned coils. For the large TRIUMF cyclotron [71], coils are separated into six different segments as the outer diameter of the coils is almost 20 m, making the transport of fully assembled coils essentially impossible over conventional roadways. The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
@@
Figure 68: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
Figure 69: Front view of the coil package in the access tunnel. The ramp is nominally engineered to a 5 m width, and maximum height of 5 m. Dimensions are in mm and degrees.
@@ 6.2 Assembly underground
The cyclotron will be assembled in a hall measuring 12 m wide, 17 m deep and 10 m high. This hall was excavated to our specifications at the 90 bend in the access ramp to the LSC, as seen in Fig. 66. A flat area, roughly $10\times 5$ m at the corner will be the staging area for components transported from the surface. A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point. Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A. [72] gave us assurances that a 40 t capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig. 70. It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
Figure 70: Examples of GH gantry cranes, capacities: foreground 20 t, background 260 tons. From Ref. [72].
Figure 71: Cyclotron and crane on cavern floor.
Figure 72: Cyclotron and crane inside the cavern.
An overview of the cyclotron inside the cavern with the rigging equipment is shown in Figures 72 - 73. It should be noted that the gantry crane will be dismantled once the heavy parts are rigged. This would increase the clearance above the cyclotron for the additional equipment. In addition, the large crane can be reconfigured and reassembled in the target hall, and used for assembly of the MEBT, target and shielding, and for subsequent maintenance of the target systems. Note, the crane is not needed for all routine maintenance of the cyclotron, as the top half is raised on jacks, providing access to the central plane and the injection, RF and extraction components.
Figure 73: Cyclotron open and crane with cavern mesh - cut-view.
6.2.1 Risks and Mitigation
Risk: Coil Size. The cyclotron coils may not fit through all of the constrictions of the mining ramp, even with the most flexible conveyance device, because of irregularities in the ramp shape and size.
Mitigation: In the case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above. Both come with drawbacks or additional cost, but are manageable.
7 Conclusion